OpenStudy (meilendurcer):

One leg of a right triangle measures 15 m and the hypotenuse measures 17 m. What is the measure of the other leg?

OpenStudy (anonymous):

\[a^2+b^2=c^2\] \[a^2+15^2=17^2\] \[a^2=17^2-15^2\] \[a=\sqrt{17^2-15^2}\]

OpenStudy (meilendurcer):

i don't follow

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

if you have a right triangle with sides \(a\) and \(b\) and hypotenuse \(c\) then pythagoras tells you that \(a^2+b^2=c^2\)

OpenStudy (anonymous):

you are given one of the legs, and the hypotenuse the hypotenuse is 17, and one leg is 15, you are asked for the other leg

OpenStudy (anonymous):

if we call the length of that leg \(a\) then by pythagoras \[a^2+15^2=17^2\] and from this you can find \(a\)

OpenStudy (meilendurcer):

64?

OpenStudy (anonymous):

yes, you get \(a^2=64\) and you want \(a\)

OpenStudy (anonymous):

since \(a^2=64\) you know that \(a=8\) since \(8^2=64\)

OpenStudy (meilendurcer):

ok so far i get it

OpenStudy (anonymous):

good, because now you are done

OpenStudy (anonymous):

measure of the other leg is \(8\)

OpenStudy (meilendurcer):

the only answer i have is 9 which is close to it

OpenStudy (anonymous):

9 is not right, 8 is right

OpenStudy (meilendurcer):

i don't have 8 :/

OpenStudy (meilendurcer):

oh oh wait sorry... i messed up thank you i do have 8

OpenStudy (anonymous):

whew

OpenStudy (meilendurcer):

yeah sorry If the hypotenuse is 41 inches, and one of the legs is 40 inches, what is the length of the other leg in inches?